OpenStudy (anonymous):

what are the first 15 counting numbers in base 9

OpenStudy (cwrw238):

as its base 9 you will have 9 digits 0 -8 you start at 1 of course - what number comes after 8?

OpenStudy (anonymous):

How do you find the first 15 counting numbers

OpenStudy (cwrw238):

1,2,3,4,5,6,7,8 - then what?

OpenStudy (cwrw238):

the next number is 10 then 11,12,13,14,15,15

OpenStudy (cwrw238):

* last one is 16

OpenStudy (anonymous):

Slightly different way of approaching this: \[1_{10}=1(10^0)\\ 1_9=1=1(9^0)\] \[2_{10}=2=2(10^0)\\ 2_9=2=2(9^0)\] ... \[8_{10}=8=8(10^0)\\ 8_9=8=8(9^0)\] \[9_{10}=9=9(10^0)\\ 9_9=10=1(9^1)+0(9^0)\] \[10_{10}=10=1(10^1)+0(10^0)\\ 10_9=11=1(9^1)+1(9^0)\] \[11_{10}=11=1(10^1)+1(10^0)\\ 11_9=12=1(9^1)+2(9^0)\] and so on.